OpenStudy (anonymous):

When you write 1,000,000 as a power of 10: the base is 1 and the exponent is 7. the base is 1 and the exponent is 6. the base is 10 and the exponent is 7. the base is 10 and the exponent is 6.

OpenStudy (anonymous):

the base is 10 and the exponent is 6.

OpenStudy (anonymous):

are you sure about that

OpenStudy (anonymous):

@Saeeddiscover is right. count the no of zeros , here its 6 place those in power of 10 which is the answer

OpenStudy (anonymous):

if the base were 1 you'd never get a number other than 1. try it! 1^2 = ? 1^5 = ? 1^(-3) = ?

OpenStudy (doc.brown):

\[10^0=1\]\[10^1=10\]\[10^2=10\times 10=100\]\[10^3=10\times\ 10\times 10=1000\]\[10^4=10\times 10\times 10\times 10=10000\]\[10^5=10\times 10\times 10\times 10\times 10=100000\]

OpenStudy (doc.brown):

\[10^0=1\]\[10^{-1}=\frac{1}{10}\]\[10^{-2}=\frac{1}{10}\times\frac{1}{10}=\frac{1}{100}\]\[10^{-3}=\frac{1}{10}\times\frac{1}{10}\times\frac{1}{10}=\frac{1}{1000}\]\[10^{-4}=\frac{1}{10}\times\frac{1}{10}\times\frac{1}{10}\times\frac{1}{10}=\frac{1}{10000}\]\[10^{-5}=\frac{1}{10}\times\frac{1}{10}\times\frac{1}{10}\times\frac{1}{10}\times\frac{1}{10}=\frac{1}{100000}\]

OpenStudy (doc.brown):

\[10^3\times 10^5=(10\times 10\times 10)\times(10\times 10\times 10\times 10\times 10)\]\[10^3\times 10^5=10\times 10\times 10\times 10\times 10\times 10\times 10\times 10\]\[10^3\times 10^5=100000000\]\[10^3\times 10^5=10^8\]

OpenStudy (doc.brown):

See how 10 to the 3+5=8?